OpenStudy (anonymous):

the rectangular coordinates are given, find polar coordinates in (r,theta) of each point (-4,4 sqrt3) a.use the equation r^2=x^2+y^2 to find the radius of the point b.use the fact that tantheta=y/x, x is not equal to 0 to find the angle (in radians) for this point b.combine your answer to write polar coordinates of this point

OpenStudy (anonymous):

x=-4 y=4 \[r^2=x^2+y^2=16+16=\]

OpenStudy (anonymous):

Can you continue from there?

OpenStudy (anonymous):

no